Determine whether g(x)=3x-6 and f(x)=1/3x+2 are inverse functions.
Thank you

Answer by fractalier(6550)

If they are, then g(f(x)) = x...let's see
g(f(x)) = g(1/3x + 2) = 3(1/3x + 2) - 6 = x + 6 - 6 = x
Yes! They are inverses!